OpenStudy (anonymous):

If U = {positive integers} and subset M = {positive odd integers}, what is M'? {positive integers} { } {1, 3, 5, 7, ...} {2, 4, 6, 8, ...}

OpenStudy (accessdenied):

Did you have any ideas on this problem so far that you weren't sure about?

OpenStudy (mayankdevnani):



OpenStudy (mayankdevnani):

M' ---- it means that the set which is complemantary to M. Or the set which is include in U(Universal Set) but not include in M(positive integers) Mathematically, \[\large \bf M'=U-M\]

OpenStudy (mayankdevnani):

so what you get ??? @kaylamarie081

OpenStudy (mayankdevnani):

U=positive integers M=positive odd integers. So, M'=U-M M'=(positive integers)-(positive odd integers) therefore,we get M'=positive even integers i.e M'={2,4,6,8......}

OpenStudy (mayankdevnani):

hope you understand. @kaylamarie081